Sol Summary – February 12th

Crew 221 Sol Summary Report 12Feb-2020


Summary Title: Robot vs Human

Author’s name: Maria Grulich, M.Sc.

Mission Status: Nominal

Sol Activity Summary:

Update since last Sol summary: The crew finally managed to have one games night which made everyone happy. Still in the wish list – movie night and dance night.

Today we reconvened for breakfast, enjoying good conversations before preparing for our only EVA of the day!

Crew commander M. Grulich, Crew Journalist H. Kalucha and Crew Greenhab Officer J. Todd went out for an EVA to install the antenna of the crew engineer.

The antenna was installed in two different places without an issue showing that such an antenna could be deployed on Mars! On the way back, the EVA team lost Fuzz, our mascot, and had to go back and rescue him! He feels like a real team member right now and the crew was scared for a second and was glad when he was found again. The team also recorded some shoots for a dance video on the EVA that was also continued in the hab: stay tuned!

After the EVA the team took a little break before heading into the robot project which needed an operation on its heart, since the raspberry pi seemed to not have been configured appropriately. The robot had to be disassembled partially to reach the LAN cable which tested the patience of Crew Journalist, the crew commander and Crew Greenhab officer. After almost giving up and letting the robot win, Crew Engineer L. Monge came to the rescue with his external monitor and keyboard as the issue was that the country and wifi was not configured.

Finally, the team saw some progress. The second batch of the microbiology samples also showed sign of microbes.

Tonight, we will have the last cultural night, Germany, and the crew commander prepared “Knödel mit Soße” and brought the most famous drink of Germany… “Apfelsaftschorle” which is apple juice with gas, a much-loved drink in Germany besides beer! Also, the crew will experience the Cologne Catherdal with VR glasses.

Anomalies in work: None

Update mouse in hab: no new mouse caught

Weather: sunny light winds

Crew Physical Status: Crew with knee issues is feeling better, some crew members experience worsened sore throats

Reports to be filed:

Sol Summary

Photo Report

Operations Report

Journalist Report

GreenHab Report

EVA Request

EVA Report 11

Journalist Report

Support Requested: None

GreenHab Report – February 12th

Crew 221 GreenHab Report 12FEB2020

GreenHab Officer: Jessica Todd

Environmental control: Heating. GreenHab door shade cloth (40%) (30%) on. Door opened today from 1:20 pm to 5:20 pm.

Average temperatures: 26°C

Hours of supplemental light: 19:30 to 00:00

Daily water usage for crops: 8.98 gallons

Daily water usage for research and/or other purposes: 0.53 gallon

Water in Blue Tank: 169.3 gallons

Time(s) of watering for crops: 1

Changes to crops: We planted some fresh lettuce, chamomile and mint! A new cucumber is growing on one of the plants. Also, the avocado seed has sprouted but there’s a prickly leafy plant also growing in the pot. Do I have permission to remove it? It’s definitely not avocado.

Narrative: Today we planted some more lettuce, since we’ve harvested quite a bit this mission. We also planted some chamomile and mint so the next crew can make some lovely fresh tea! Our corn is growing nicely, we now have corn growing out of purely Martian soil. Unfortunately no progress for our mungbeans, apparently they don’t like Mars.

Harvest: 21g Rosemary, 26g Butter lettuce, 2g Oregano, 1g Basil, 2g Chives.

Support/supplies needed: NA

Other Notes: Still a bit worried about the water level in the GreenHab tank (see previous GreenHab report). Please advise.

Journalist Report – February 12th

Crew 221 Journalist Report 12Feb2020

Author: Hemani Kalucha, Crew Journalist

Sol 10

Title: Hilarious Heads Up

This morning, the all girl (Maria, Hemani, Jess) EVA team headed out to complete radio assembly. The team successfully built the radio tower without help from HabCom in two different locations. After the mission objective was complete, the team made some pit stops on the way back to film their dance video, set to the music of “We Are Not Alone” from the move, the Breakfast Club. This will be used for outreach. The music fit perfectly, because we are trying to survive as life on Mars, and we are in a remarkably similar situation to the characters of the breakfast club – six strangers getting to know each other in a confined space.

After a yummy lunch, the girls got back into programming the robot, and with the help of Crew Engineer Luis, solved the software issue! This was such encouraging news, and we are hoping to test it out in the field tomorrow! Then GHO Jess and Hemani planted new crops in the Greenhab – Chamomile, Spearmint, and Lettuce. Before dinner, the crew played a little Heads Up, which lead to some hilarious charades. We are about to sit down to a dinner of delicious German cuisine and cultural night by our very own Commander!

Operations Report – February 12th

Crew 221 Operations Report 12Feb2020

SOL: 10

Name of person filing report: Luis D. Monge

Non-nominal systems:

Notes on non-nominal systems:

Generator: on (at 7:40 pm)

Hours run: 13:00

From what time last night: 7:40pm

To what time this morning: 7:00am

List any additional daytime hours when the generator was run : 0

Solar— SOC % (Before generator is run at night) 79%

Diesel Reading – 85%

Station Propane Reading – 50%

Ethanol Free Gasoline – N/A

Water (loft tank): 35 gallons

Water Meter: 148 units

Water (static tank) – 215 gals

Static to Loft Pump used – no

Water in GreenHab – 169.3 gals (water is probably lower than estimated, please refer to greenhab report)

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: no


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: no

Hours: (before EVA) : 130.4

Beginning charge: (Before EVA) 100%

Ending charge: (On return from EVA, before recharging)

Currently charging: yes

Opportunity rover used: no

Hours: n/a

Beginning charge: n/a

Ending charge: n/a

Currently charging: yes

Curiosity rover used: no

Hours: 136.6

Beginning charge: 100%

Ending charge:

Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) not used

Reason for use: not used

Oil Added? not used

ATV Fuel Used: Gals not used

# Hours the ATVs were Used today: not used

Notes on ATVs:

HabCar used and why, where? not used

CrewCar used and why, where? not used

General notes and comments: not used

Summary of internet: operational

Summary of suits and radios: operational; 4 blue radios and

Summary of Hab operations: nominal.

Summary of GreenHab operations: operational, black corn experiment continues, now watered with 2l/day. watering of the plants carried out as per procedure

Summary of ScienceDome operations: one oven is operating in the lab, the incubator is non-functional, microbiology experiments going on and cultures are bieng grown in agar plates.

Summary of RAM operations: robot assembly almost completed and expected to continue, radio comunication experiment continued expected to finish tomorrow.

Summary of any observatory issues: nominal

Summary of health and safety issues: crew memeber with flu continues recovery.

Questions, concerns and requests to Mission Support: nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site